The CFO of a consulting engineering firm is deciding between purchasing Ford Explorers and Toyota 4Runners for company principals. The purchase price for the Ford Explorer will be $31,000. Annual maintenance costs for the Explorer are expected to be $600 per year more than that of the 4Runner. The purchase price for Toyota 4Runners is 35,000 The trade-in values after 3 years are estimated to be 50% of the first cost for the Explorer and 60% for the 4Runner. (a) What is the incremental ROR between the two vehicles? (b) Provided the firm’s MARR is 18% per year, which vehicle should it buy?
Answer #1

ANSWER:

In order to find the incremental ror , we will equate the incremental pw to zero.

we will subtract ford explorer from toyota 4 runners as the purchase price is more of toyota 4 runners.

salvage value of toyota 4 runners = 50% * purchase price = 50% * 31,000 = 15,500

salvage value of ford explorer = 60% * purchase price = 60% * 35,000 = 21,000

difference in salvage value = 21,000 - 15,500 = 5,500

pw (toyota 4 runners - ford explorer) = purchase price (toyota 4 runners - ford explorer) + amc (toyota 4 runners - ford explorer) (p/a,i,n) + salvage value (toyota 4 runners - ford explorer) (p/f,i,n)

0 = (-35,000 - (-31,000) ) + (-600) (p/a,i,3) + 5,500(p/f,i,3)

0 = -4,000 - 600(p/a,i,3) + 5,500(p/f,i,3)

solving via trial and error we get that i is between -2% and -3% and solving further we get that i is -2.22%

so the incremental ror is -2.22%

b) since the marr is more then incremental irr , therefore it should buy the toyota 4 runners.